A flexible rectangular coil of wire with 150 loops is stretched in such a way that its dimension changes from 5 cm x 8 cm to 7 cm x 11 cm in 0.15 s in a magnetic field of 25T that is 30 degrees relative to the plane of the coil.

a. Calculate the induced EMF


Show full solution:


Given:


Solution:


1
Expert's answer
2022-05-24T13:56:37-0400

"\varepsilon=\frac{n∆SB\cos\alpha}t,"

"\u2206S=S_2-S_1=cd-ab,"

"\u03b5= \\frac{\n\nn(cd-ab)Bcos\u03b1}t\n\u200b\n =46.3 ~V."